OpenStudy (anonymous):

Simplify the expression below. -5-8

OpenStudy (nory):

-5 - 8 = -13 Think of it this way: 5 + 8 = 13, but since it's negative, it's -5 - 8 = -13

OpenStudy (anonymous):

subtract 8 from -5 -13

OpenStudy (anonymous):

Suppose you owe 5 dollars and you borrow another 8, how much would you owe then?

OpenStudy (nory):

@domu I like that way of thinking about it.

OpenStudy (anonymous):

: )

OpenStudy (anonymous):

I always use money to solve my problems haha